### v1

# Enhancing Person-to-Person Virtual Try-On with Multi-Garment Virtual Try-Off

Abstract
Computer vision is transforming fashion through Virtual Try-On (VTON) and Virtual Try-Off (VTOFF). VTON generates images of a person in a specified garment using a target photo and a standardized garment image, while a more challenging variant, Person-to-Person Virtual Try-On (p2p-VTON), uses a photo of another person wearing the garment. VTOFF, on the other hand, extracts standardized garment images from clothed individuals. We introduce TryOffDiff, a diffusion-based VTOFF model. Built on a latent diffusion framework with SigLIP image conditioning, it effectively captures garment properties like texture, shape, and patterns. TryOffDiff achieves state-of-the-art results on VITON-HD and strong performance on DressCode dataset, covering upper-body, lower-body, and dresses. Enhanced with class-specific embeddings, it pioneers multi-garment VTOFF, the first of its kind. When paired with VTON models, it improves p2p-VTON by minimizing unwanted attribute transfer, such as skin color. Code is available at: https://rizavelioglu.github.io/tryoffdiff/
Figure 1: Virtual try-off results generated by multi-garment TryOffDiff. Rows display the input reference images, model predictions, and the ground truth. Conditioned on the person image and a specified garment class, TryOffDiff reconstructs: upper-body garments VITON-HD (cols 1-3), Dresscode (cols 4-6); lower-body garments (cols 7-9) and dresses (cols 10-12).
1 Introduction
Image-based virtual try-on (VTON) [19] is a key computer vision task aimed at generating images of a person wearing a specified garment. Traditionally, two input images are required: one showing the garment in a standardized form (often from an e-commerce catalog) and another of the target person that needs to be ‘dressed’ [8, 22]. Recent methods focus on a modified formulation where the catalog image is replaced with a photo of another person wearing the target garment, the so-called person-to-person try-on (p2p-VTON). This introduces additional processing complexity [38] as the model does not have access to full garment information.
From an application perspective, VTON offers an interactive shopping experience that helps users make better-informed purchasing decisions. On the research side, it raises intriguing research questions, particularly around human pose detection as well as clothing shape, pattern, and texture analysis [15]. Best-performing models are usually guided generative models focused on creating specific, physically accurate outputs. Unlike general generative tasks that produce diverse outputs, reconstruction requires models to generate images that align with the correct appearance of the garment on a person. However, one major drawback of VTON is the lack of a clearly defined target output, often resulting in stylistic variations that complicate consistent evaluation. As an example, generated images may show garments tucked, untucked, or altered in fit, introducing plausible yet inconsistent visual variations and make it difficult to assess the true quality of the model.
A related task, Virtual Try-OFF (VTOFF) [34], focuses on the reverse goal: generating standardized garment images from photos of clothed persons. Despite sharing challenges with VTON—such as pose analysis, geometric transformations, and preserving textures, patterns, and logos—VTOFF offers a key advantage: its outputs have lower stylistic variability (e.g., no tucked vs. untucked variations), making reconstruction quality easier to assess. However, this line of research remains scarce [37, 32]. From an application perspective, VTOFF can enhance e-commerce by providing consistent garment images for integration into VTON systems, such as OOTDiffusion [40], enabling practical p2p-VTON without ground truth garment images.
In this paper we introduce TryOffDiff, a diffusion-based model tailored to the specific challenges of VTOFF task. The base variant mirrors Stable Diffusion architecture and replaces text conditioning with image conditioning, using SigLIP features [43] processed by an adapter module to capture garment-specific properties such as texture, shape and patterns. Our approach achieves state-of-the-art results on the VITON-HD dataset for upper-body garment reconstruction. To support multi-garment scenarios, we extend TryOffDiff by adding class-specific embeddings to timestep embeddings [24]. We demonstrate our results on the full DressCode dataset and show that the multi-garment TryOffDiff achieves comparable results to garment-specific variants. This makes TryOffDiff the first model to tackle multi-garment virtual try-off. Finally, we integrate TryOffDiff with state-of-the-art VTON models to enable person-to-person try-on. Our results show that this approach outperforms models capable of p2p-VTON by reducing unwanted attribute transfer, such as skin color, from the source person. To summarize, our primary contributions are:
  • TryOffDiff: A diffusion model tuned for virtual try-off, excelling at garment reconstruction.
  • Multi-Garment VTOFF: The first approach capable of handling multiple garment categories, and enabling simultaneous reconstruction of multiple garments, achieved through class-specific embeddings.
  • Enhanced Person-to-Person VTON: Improved performance by feeding TryOffDiff output into VTON systems.
Figure 2: Overview of various fashion image generation pipelines. First row illustrates Virtual Try-On, which synthesizes an image of a person wearing a target garment given input images of person and garment. Second row depicts Virtual Try-Off, which generates an e-commerce-style image of a garment from a photo of a person wearing it. Third row shows Person-to-Person Virtual Try-On, where a garment is transferred from one person to another. Finally, fourth row demonstrates that P2P-VTON can be achieved by integrating VTON and VTOFF models.
2 Related Work
This section reviews prior research relevant to our TryOffDiff model, covering Virtual Try-Off, Image-based Virtual Try-On, and Conditional Diffusion Models. Virtual Try-Off generates standardized garment images from clothed individuals, complementing virtual fitting tasks. Image-based Virtual Try-On synthesizes personalized garment images using catalog or person-worn inputs, supporting advanced variants. Conditional Diffusion Models offer the generative backbone for our approach, enabling controlled synthesis through tailored conditioning mechanisms. These areas highlight the challenges and strategies shaping our work on multi-garment reconstruction.
Virtual Try-Off.
Virtual Try-Off (VTOFF) generates standardized garment images from photos of clothed persons, a task with growing interest due to its applications in e-commerce. Early works, such as TileGAN [42], used a two-stage approach: a U-Net-inspired encoder-decoder to transfer features like color and shape for a coarse image, followed by a pix2pix-based refinement step. With the rise of general-purpose text-to-image diffusion models like Stable Diffusion [28], later works shifted toward text-driven garment generation. ARMANI [47], DiffCloth [48], GarmentAligner [46], and MGD [2] explored finetuning latent diffusion models on garment-caption datasets. However, text-based approaches require highly detailed captions, which are difficult to produce, and often struggle to follow instructions precisely, leading to suboptimal outputs.
To address these limitations, recent approaches adopted image-based conditioning. IGR [31] finetuned Stable Diffusion v1.5 with a dual-tower architecture. Similarly, RAGDiffusion [32] employed SDXL [26] with a dual-tower design and a coarse mask from a complex retrieval-augmented generation. TryOffAnyone [37] finetuned Stable Diffusion v1.5 for VTOFF, training only self-attention layers (similar to CatVTON [10]), without additional image encoders. This resulted in a lightweight yet powerful model, making it a strong baseline. Despite these advances, VTOFF research faces challenges: many studies are not publicly available, limiting open exploration, and most focus solely on upper-body garments, leaving broader garment categories underexplored.
Image-based Virtual Try-On.
The objective of image-based virtual try-on is to create composite images of a target person wearing a specific garment, preserving the person’s identity, pose, and body shape while accurately rendering garment details. CAGAN [19] introduced this task with a cycle-GAN approach, while VITON [15] formalized it as a two-step, supervised framework: warping the garment through non-parametric geometric transformations [4], followed by blending it onto the person. CP-VTON [35] improved warping with a learnable thin-plate spline (TPS) transformation via a geometric matcher, later enhanced by dense flow [16] and appearance flow [14] for better pixel-level alignment of textures and folds. Despite these advanves, warping-based methods struggled with complex garment features like logos and intricate patterns.
Recent efforts shifted to GAN-based and diffusion-based models to overcome these limitations. FW-GAN [12] synthesized try-on videos, while PASTA-GAN [38] adapted StyleGAN2 for person-to-person try-on. However, GANs faced challenges with unstable training and mode collapse, prompting a move toward diffusion models, which offer greater stability. M&M-VTO [49] introduced a single-stage diffusion model for multi-garment try-on using one person image and multiple garment inputs. IDM-VTON [9] used two modules to encode garment semantics, extracting high- and low-level features with cross-attention and self-attention layers. OOTDiffusion [40] leveraged pretrained latent diffusion models, integrating garment features into a denoising UNet via outfitting fusion. In a lighter approach, CatVTON [10] avoided heavy feature extraction, proposing a compact model from a pretrained latent diffusion model with promising results using fewer parameters.
Adapting VTON models for VTOFF is challenging, as VTON often relies on additional inputs like text prompts, keypoints, or segmentation masks, requiring careful adjustment. Importantly, while both VTON and VTOFF tasks involve garment manipulation, they differ fundamentally. VTON models have access to complete garment details, allowing them to primarily focus on warping the item to fit a target pose. In contrast, VTOFF models must work with only partial garment information from a reference image, where occlusions and deformations are common, requiring them to reconstruct missing details from limited visual cues.
Conditional Diffusion Models.
Latent Diffusion Models [28] (LDMs) achieved great success in recent years, providing control over generative process through cross-attention mechanisms [33]. These models support conditioning on various input modalities, including text [5, 3, 13] and images [29, 25, 30]. In text-guided synthesis, ControlNet [44] and T2I-Adapter [23] enhance pretrained models with additional layers for precise spatial control. IP-Adapter [41] further improves flexibility by decoupling cross-attention for text and image features, allowing image-guided generation with optional structural conditions. Prompt-Free Diffusion [39] eliminates text prompts entirely, relying on reference image and optional structural inputs to guide output.
Despite progress, applying these models to garment reconstruction poses challenges. Text-guided methods require highly detailed prompts for each sample to define product attributes, which is impractical. Similarly, existing image-guided models lack the ability to enforce the strict standards of fashion photography [7], limiting their suitability.
3 Methodology
Figure 3: Overview of TryOffDiff. Given a reference image and a class label (e.g. ‘dresses’), the SigLIP [43] image encoder extracts features, which are subsequently processed by a learnable adapter and embedded into the cross-attention layers of the Denoising U-Net. A learnable class label embedding conditions the generation to support multi-garment reconstruction. The model, with trainable Adapter, Embedding, and U-Net components, produces a standardized garment image, which is decoded by a frozen VAE.
This section formally defines the virtual try-off task and outlines TryOffDiff model, including SigLIP-based image conditioning and class-conditioning for multi-garment support.
3.1 Virtual Try-Off
Problem Formulation.
Consider an RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ with height $H\in\mathbb{N}$ and width $W\in\mathbb{N}$, representing a person wearing garments. The VTOFF task is to produce a standardized product image $\mathbf{G}\in\{0,\ldots,255\}^{H\times W\times 3}$, that meets commercial catalog standards. Formally, the goal is to train a generative model that learns the conditional distribution $P(G|C)$, where $G$ and $C$ represent the variables corresponding to garment images and reference images (serving as condition), respectively. Suppose the model approximates this target distribution with $Q(G|C)$. Then, given a specific reference image $\mathbf{I}$ as conditioning input, the objective is for a sample $\hat{\mathbf{G}}\sim Q(G|C=\mathbf{I})$ to resemble a true sample of a garment image $\mathbf{G}\sim P(G|C=\mathbf{I})$ as closely as possible.
3.2 TryOffDiff
The architecture of TryOffDiff mirrors Stable Diffusion [28] (v1.4), a latent diffusion model originally designed for text-to-image generation using CLIP’s [27] text encoder. We replace text prompts for direct image-guided generation.
Image Conditioning.
A core challenge in image-guided generation is effectively incorporating visual features into the conditioning mechanism of the generative model. CLIP’s ViT [27] has become a popular choice for image feature extraction due to its general-purpose capabilities. Recently, SigLIP [43] introduced modifications that improve performance, particularly for tasks requiring more detailed and domain-specific visual representations. Therefore, we use the SigLIP model as image feature extractor and retain the full sequence of token representations in its final layer to preserve spatial information essential for accurate garment reconstruction. These features are processed by a trainable adapter module:
$\mathbf{C}(\mathbf{I})=(\mathrm{LN}\circ\mathrm{Linear}\circ\mathrm{SigLIP})(\mathbf{I})\in\mathbb{R}^{n\times m}$ (1)
where SigLIP embeddings are projected by a linear layer and normalized (LN) [1], following a structure similar to IP-Adapter [41]. The adapted features $\mathbf{C}(\mathbf{I})$ are then incorporated into the denoising U-Net’s cross-attention layers. Specifically, the key $\mathbf{K}$ and value $\mathbf{V}$ of the attention mechanism at each layer are derived from the image features through linear transformations:
$\mathbf{K}=\mathbf{C}(\mathbf{I})\cdot\mathbf{W}_{k}\in\mathbb{R}^{n\times d_{k}},\mathbf{V}=\mathbf{C}(\mathbf{I})\cdot\mathbf{W}_{V}\in\mathbb{R}^{n\times d_{v}}$ (2)
where $\mathbf{W}_{k}\in\mathbb{R}^{m\times d_{k}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{m\times d_{v}}$. This mechanism allows the U-Net to use the reference image’s visual details, improving the precision of the generated garment.
[…]
TryOffAnyone* [37] - - - 72.0 - - 17.2 25.3 - 8.9 21.0
TryOffDiff-single CLIP ViT-B/32 - (257,1024) - 77.9 69.2 46.4 35.3 16.8 7.7 5.2 23.7
CLIP ViT-B/32 Linear+LN (77,768) - 78.8 70.5 46.5 34.3 16.3 7.4 4.7 23.5
SigLIP-B/16 Linear+LN (77,768) - 78.4 69.8 47.3 33.1 16.4 6.6 5.1 22.1
CLIP ViT-B/32 Linear+LN (77,768) 79.8 71.5 47.8 32.0 15.2 5.8 4.1 21.5
SigLIP-B/16 Linear+LN (77,768) 80.3 72.5 49.2 30.1 14.5 5.7 3.9 20.3
Table 1: Training configurations and quantitative results of baseline and ablation studies on VITON-HD-test. The “Pretrain” check mark indicates whether the U-Net is initialized from Stable Diffusion v1.4 (✓) or trained from scratch (-). All ablations mirror the model architecture of Stable Diffusion (v1.4). *Values for TryOffAnyone are taken from the original paper.
Garment Type Conditioning.
Reference images often contain multiple garments, such as a shirt and pants, requiring the model to distinguish which garment to reconstruct. To resolve this ambiguity and enable multi-garment support, we condition the model on text embeddings corresponding to garment categories—“lower body”, “upper body”, or “dress”—aligned with the class labels of the Dress Code dataset. These text embeddings are passed through a linear layer and then combined with timestep embeddings [24] via element-wise addition before being fed into the residual blocks of the diffusion model. This conditioning mechanism guides the denoising process to generate the garment corresponding to specified category.
4 Experiments
Since VTOFF task is newly introduced, few open-source methods exist. We evaluate our proposed TryOffDiff model against the publicly available TryOffAnyone [37] to establish a baseline. To ensure reproducibility, we provide a detailed description of our experimental setup. We adopt DISTS [11] as the primary evaluation metric, supplemented by standard metrics for a comprehensive comparison. Qualitative results are also presented to showcase TryOffDiff’s ability to handle diverse and challenging inputs.
4.1 Experimental Setup
Datasets.
Our experiments are conducted on two publicly available datasets: VITON-HD [8] and Dress Code [22]. Originally designed for VTON, these datasets are also well-suited for VTOFF as they provide image pairs $(\mathbf{I},\mathbf{G})$, where $\mathbf{I}$ is a clothed person and $\mathbf{G}$ is the corresponding garment.
VITON-HD includes $13,679$ high-resolution ($1024\times 768$) image pairs of frontal half-body models with upper-body garments. We preprocessed the dataset by removing duplicates and test set leaks from the training set, resulting in $11,552$ unique pairs for training and $1,990$ for testing.
Dress Code contains $53,792$ high-resolution ($1024\times 768$) person-garment pairs, with $48,392$ pairs for training–comprising dresses ($27,678$), upper-body ($13,563$), and lower-body ($7,151$)–and $5,400$ pairs for testing, evenly split with $1,800$ pairs per category.
For person-to-person try-on, where ground-truth images are unavailable, we randomly pair garments across individuals and provide a text file of input image filenames for reproducibility. For Dress Code, a label (upper-body, lower-body, or dresses) indicates the garment to transfer; this is not needed for VITON-HD, which includes only upper-body garments.
Evaluation Metrics.
To evaluate reconstruction quality, we use established full-reference metrics such as Structural Similarity Index Measure (SSIM) [36] and its multi-scale (MS-SSIM) as well as complex-wavelet (CW-SSIM) variants. To measure perceptual quality, we use common no-reference metrics like Learned Perceptual Image Patch Similarity (LPIPS) [45], Fréchet Inception Distance (FID) [17] and Kernel Inception Distance (KID) [6] metric. To measure the perceptual similarity between images capturing both structural and textural information, we report the Deep Image Structure and Texture Similarity (DISTS) [11] metric. We prioritize the DISTS metric in evaluations because it balances structural and textural information, offering a comprehensive evaluation of image quality.
Implementation Details.
Denoising U-Net weights are initialized from Stable Diffusion v1.4 [28] and subsequently fine-tuned. The Adapter module and class Embedding layer are trained from scratch. The SigLIP encoder, VAE encoder, and VAE decoder are kept frozen throughout the training, which is performed end-to-end. Input reference images are padded along the width to achieve a square aspect ratio and resized to $512\times 512$ to match the pretrained SigLIP and VAE encoder’s format. Garment images undergo the same preprocessing during training. We use SigLIP-B/16-512 as image feature extractor, yielding 1,024 token embeddings of dimension 768, which the adapter–comprising a linear layer and normalization–reduces to $n=77$ conditioning embeddings of dimension $m=768$. For garment conditioning, we embed class labels (e.g. ‘upper-body’, ‘lower-body’, ‘dress’) using a trainable embedding layer. This layer has 3 classes and projects the labels into a 1,280-dimensional space, matching the dimension of the timestep embeddings. These class embeddings are then combined with the timestep embeddings via element-wise addition before being fed into the U-Net’s residual blocks.
Training runs for 200 epochs (approximately 150k iterations) on a single node with four NVIDIA A40 GPUs, taking about 5 days with a batch size of 16. We use the AdamW optimizer [21], with an initial learning rate of 5e-5, which increases linearly from zero over the first 15,000 warmup steps (%10 of total iterations), then decays linearly to zero. A weight decay of 0.01 applies to all parameters except biases and normalization weights. Following [20], we use the Euler noise scheduler (Algorithm 2) with 1,000 steps. Optimization relies on the standard Mean Squared Error (MSE) loss, which measures the distance between the added and the predicted noise at each step [18].
During inference, TryOffDiff uses Euler scheduler with 20 timesteps and a guidance scale of 1.5. On a single NVIDIA A40 GPU, inference on Dress Code test set takes 2.8 seconds per image and requires 10.1GB of memory.
4.2 Ablation Studies
We perform ablation studies to assess key components of TryOffDiff, including the image encoder, adapter module, conditioning feature dimensions, and the impact of finetuning a pretrained U-Net versus training from scratch. The configurations for these experiments are summarized in Tab. 1. In addition, Fig. 4 illustrates the impact of guidance scale and inference steps on FID and DISTS scores, evaluated on the Dress Code-test set using the Euler noise scheduler. This experiment reveals how these hyperparameters influence reconstruction quality, providing insights into optimal inference setting for TryOffDiff.
(a) Guidance Scale
(b) Inference steps
Figure 4: Impact of guidance scale ($s$) and inference steps ($n$) on DISTS and FID scores. Experiments are conducted on DressCode-test with TryOffDiff using the Euler scheduler [20].
4.3 Quantitative Results
We evaluate TryOffDiff’s performance on the VITON-HD and Dress Code datasets, comparing it to baseline TryOffAnyone [37] and category-specific ablations. Results are presented in Tab. 1 (VITON-HD), Tab. 2 (Dress Code), and Tab. 3 (p2p-VTON).
[…]
Dresses
TryOffDiff-single 81.6 76.1 55.6 26.1 18.4 8.2 4.4 20.8
TryOffDiff-multi 81.6 76.1 55.5 26.4 18.9 8.0 4.6 21.2
Upper-body
TryOffDiff-single 80.8 73.8 47.8 31.6 17.1 5.2 4.7 21.6
TryOffDiff-multi 80.2 73.2 48.1 32.3 17.8 5.3 5.5 22.2
Lower-body
TryOffDiff-single 81.1 76.8 55.5 30.0 22.6 9.4 5.9 21.1
TryOffDiff-multi 80.4 75.6 54.1 30.8 22.4 9.0 6.8 21.7
All
TryOffDiff-merged 81.2 75.6 52.9 29.2 11.4 6.7 3.1 21.2
TryOffDiff-multi 78.6 75.4 57.4 25.0 12.8 6.8 4.1 21.3
Table 2: Quantitative comparison of multi-garment VTOFF on DressCode-test. We compare category-specific models (TryOffDiff-single) trained independently on dresses, upper-body, and lower-body garments against a class-conditioned model (TryOffDiff-multi) trained jointly across all garment types. Additionally, we report results for TryOffDiff-merged, which aggregates the outputs of the three category-specific models into a single set. While non-distributional metrics (e.g., SSIM, LPIPS) for merged results reflect simple averages across categories, distributional metrics (e.g., FID, KID) are computed over the merged output set, capturing holistic distributional properties.
VITON-HD Dress Code
[…]
CatVTON 12.0 3.5 3.9 8.4 1.9 3.1
OOTD + GT 10.8 2.8 2.0 7.5 3.4 2.5
OOTD + TryOffDiff 11.9 3.3 2.6 7.9 3.4 2.7
Table 3: Quantitative results for p2p-VTON. We evaluate OOTDiffusion (OOTD) with ground truth (GT) garments and TryOffDiff-predicted garments, alongside CatVTON, a specialized p2p-VTON model. Our TryOffDiff model, when integrated with a VTON model, achieves competitive performance, even though it is not explicitly trained for the p2p task.
The results in Tab. 1 reveal key insights into our design choices. First, finetuning a pretrained U-Net consistently outperforms training from scratch across all metrics, underscoring the advantage of leveraging pretrained weights for initializing the diffusion model. Second, the SigLIP encoder outperforms CLIP, especially in perceptual quality metrics like FID and DISTS, suggesting that SigLIP’s visual representations are more effective for detailed garment reconstruction. Finally, incorporating the adapter module further improves performance by more effectively conditioning the diffusion process with image features.
On VITON-HD dataset (Tab. 1), our method outperforms the baseline across all metrics except LPIPS, where TryOffAnyone achieves a better score, though TryOffDiff excels in global metrics like FID and DISTS which is our primary metric.
In our evaluation on the Dress Code dataset (Tab. 2), we consider: (i) category-specific TryOffDiff models trained individually on dresses, upper-body, and lower-body garments; (ii) a class-conditioned TryOffDiff model trained jointly on all categories using garment-type embeddings; and (iii) an aggregated result, denoted TryOffDiff-merged, where outputs from the three category-specific models are combined post hoc. The class-conditioned model remains competitive across all categories, demonstrating its robustness as a unified solution.
Figure 5: Qualitative comparison of garment-specific and multi-garment TryOffDiff. First row displays the conditioning image. Second row shows dataset-specific reconstructions produced by TryOffDiff model trained on individual datasets: VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Last row shows predictions of multi-garment TryOffDiff trained on full DressCode with garment type conditioning.
For person-to-person virtual try-on (Tab. 3), we integrate TryOffDiff with OOTDiffusion and compare it to CatVTON, a specialized p2p-VTON model. On VITON-HD, OOTD+TryOffDiff performs slightly better than CatVTON. On Dress Code, TryOffDiff, when paired with OOTD, delivers competitive performance, approaching specialized p2p-VTON models like CatVTON, especially in perceptual quality metrics.
These findings highlight TryOffDiff’s versatility and effectiveness, enabling high-quality garment reconstruction for both VTOFF and p2p-VTON tasks. Its improvements in structural similarity and perceptual quality translate into practical benefits, such as more realistic garment rendering in both VTOFF and p2p-VTON applications.
4.4 Qualitative Analysis
Figure 5 illustrates the performance of TryOffDiff in garment reconstruction, comparing two variants: garment-specific models (row 2) and multi-garment model (row 3). The conditioning images, displayed in first row, are sourced from VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12). Garment-specific models excel in precision, particularly for text and fine details (cols 2 and 5), likely due to the pretrained model initialization. In contrast, the multi-garment model offers versatility, delivering strong performance across diverse garment types and datasets, albeit with slight trade-offs in detail compared to specialized models. Nevertheless, it reconstructs a high level of detail (column 6) and shows robust generalization to VITON-HD, which was unseen during training. In some cases (cols 2, 7 and 9), it even produces better fidelity of color.
Figure 6 illustrates the results of p2p-VTON. Overall, neither method stands out as a definitive winner. CatVTON excels in preserving texture and pattern details but occasionally introduces diffusion artifacts (row 2). It can also transfer unintended attributes, such as skin color, from target model to the source model (row 1)–a limitation not observed in classical try-on models. OOTDiffusion combined with TryOffDiff avoids such attribute transfer issues but may struggle due to masking, resulting in residual source garment attributes in the output (rows 3 and 4).
These results highlight the potential of Virtual Try-Off task and the TryOffDiff model. Although TryOffDiff was not specifically trained for p2p-VTON, its integration with VTON models presents a promising approach, already demonstrating competitive performance compared to state-of-the-art person-to-person virtual try-on methods.
Figure 6: Qualitative comparison for Person-to-Person Virtual Try-On. Columns show: (a) person to be dressed, (b) CatVTON output using a person image with target garment as condition for direct p2p-VTON, (c) OOTDiffusion output with ground truth target garment, and (d) OOTDiffusion output with multi-garment TryOffDiff prediction.
5 Conclusion
In this paper we propose TryOffDiff, a latent diffusion model tailored to Virtual Try-Off. We substitute text conditioning mechanism with adapted SigLIP features to steer the generative process, enabling effective garment reconstructions. TryOffDiff achieves state-of-the art results on VITON-HD. Further experiments were conducted on DressCode, which contains three types of garments: upper-body, lower-body and dresses. While the baseline architecture can be trained separately for each category, we enhance TryOffDiff by introducing class conditioning. This advancement allows a single model to perform multi-garment reconstruction, successfully capturing all garment types with minimal loss in detail quality. To our knowledge, this is the first method to address multi-garment VTOFF in a unified framework.
Moreover, we demonstrate that integrating VTOFF with VTON models enables Person-to-Person Virtual Try-On, achieving performance comparable to specialized models. This pipeline offers practical advantages: it leverages existing VTON models for a wider applications and decouples garment reconstruction from other generative artifacts, simplifying the assessment of p2p-VTON outcomes.
Despite these promising results, challenges persist in preserving complex structures, such as logos and printed designs. Future work could explore advanced generative models, alternative visual conditioning methods and additional losses to enhance detail preservation.
Acknowledgment
This work has been funded by Horizon Europe program under grant agreement 101134447-ENFORCE, and by the German federal state of North Rhine-Westphalia as part of the research funding program KI-Starter. We would like to thank UniZG-FER for providing access to their hardware.

### v2

@@
-Enhancing Person-to-Person Virtual Try-On with Multi-Garment Virtual Try-Off
+MGT: Extending Virtual Try-Off to Multi-Garment Scenarios
 Abstract
-Computer vision is transforming fashion through Virtual Try-On (VTON) and Virtual Try-Off (VTOFF).
+Computer vision is transforming fashion industry through Virtual Try-On (VTON) and Virtual Try-Off (VTOFF).
 VTON generates images of a person in a specified garment using a target photo and a standardized garment image, while a more challenging variant, Person-to-Person Virtual Try-On (p2p-VTON), uses a photo of another person wearing the garment.
-VTOFF, on the other hand, extracts standardized garment images from clothed individuals.
-We introduce TryOffDiff, a diffusion-based VTOFF model.
-Built on a latent diffusion framework with SigLIP image conditioning, it effectively captures garment properties like texture, shape, and patterns.
-TryOffDiff achieves state-of-the-art results on VITON-HD and strong performance on DressCode dataset, covering upper-body, lower-body, and dresses.
-Enhanced with class-specific embeddings, it pioneers multi-garment VTOFF, the first of its kind.
-When paired with VTON models, it improves p2p-VTON by minimizing unwanted attribute transfer, such as skin color.
-Code is available at: https://rizavelioglu.github.io/tryoffdiff/
-Figure 1: Virtual try-off results generated by multi-garment TryOffDiff.
-Rows display the input reference images, model predictions, and the ground truth.
-Conditioned on the person image and a specified garment class, TryOffDiff reconstructs: upper-body garments VITON-HD (cols 1-3), Dresscode (cols 4-6); lower-body garments (cols 7-9) and dresses (cols 10-12).
+VTOFF, in contrast, extracts standardized garment images from photos of clothed individuals.
+We introduce Multi-Garment TryOffDiff (MGT), a diffusion-based VTOFF model capable of handling diverse garment types, including upper-body, lower-body, and dresses.
+MGT builds on a latent diffusion architecture with SigLIP-based image conditioning to capture garment characteristics such as shape, texture, and pattern.
+To address garment diversity, MGT incorporates class-specific embeddings, achieving state-of-the-art VTOFF results on VITON-HD and competitive performance on DressCode.
+When paired with VTON models, it further enhances p2p-VTON by reducing unwanted attribute transfer, such as skin tone, ensuring preservation of person-specific characteristics.
+Demo, code, and models are available at: https://rizavelioglu.github.io/tryoffdiff/
+Figure 1: Virtual try-off results generated by Multi-Garment TryOffDiff (MGT).
+The first three columns demonstrate MGT’s ability to generate multi-garment images (e.g., upper- and lower-body garments) from a single reference image.
+The last two columns illustrate dress reconstruction results.
+MGT generates an image in under 3 seconds on a consumer GPU.
+Please zoom in for a clearer view of details.
 1 Introduction
-Image-based virtual try-on (VTON) [19] is a key computer vision task aimed at generating images of a person wearing a specified garment.
-Traditionally, two input images are required: one showing the garment in a standardized form (often from an e-commerce catalog) and another of the target person that needs to be ‘dressed’ [8, 22].
-Recent methods focus on a modified formulation where the catalog image is replaced with a photo of another person wearing the target garment, the so-called person-to-person try-on (p2p-VTON).
-This introduces additional processing complexity [38] as the model does not have access to full garment information.
-From an application perspective, VTON offers an interactive shopping experience that helps users make better-informed purchasing decisions.
-On the research side, it raises intriguing research questions, particularly around human pose detection as well as clothing shape, pattern, and texture analysis [15].
-Best-performing models are usually guided generative models focused on creating specific, physically accurate outputs.
-Unlike general generative tasks that produce diverse outputs, reconstruction requires models to generate images that align with the correct appearance of the garment on a person.
-However, one major drawback of VTON is the lack of a clearly defined target output, often resulting in stylistic variations that complicate consistent evaluation.
-As an example, generated images may show garments tucked, untucked, or altered in fit, introducing plausible yet inconsistent visual variations and make it difficult to assess the true quality of the model.
-A related task, Virtual Try-OFF (VTOFF) [34], focuses on the reverse goal: generating standardized garment images from photos of clothed persons.
-Despite sharing challenges with VTON—such as pose analysis, geometric transformations, and preserving textures, patterns, and logos—VTOFF offers a key advantage: its outputs have lower stylistic variability (e.g., no tucked vs.
-untucked variations), making reconstruction quality easier to assess.
-However, this line of research remains scarce [37, 32].
-From an application perspective, VTOFF can enhance e-commerce by providing consistent garment images for integration into VTON systems, such as OOTDiffusion [40], enabling practical p2p-VTON without ground truth garment images.
-In this paper we introduce TryOffDiff, a diffusion-based model tailored to the specific challenges of VTOFF task.
-The base variant mirrors Stable Diffusion architecture and replaces text conditioning with image conditioning, using SigLIP features [43] processed by an adapter module to capture garment-specific properties such as texture, shape and patterns.
-Our approach achieves state-of-the-art results on the VITON-HD dataset for upper-body garment reconstruction.
-To support multi-garment scenarios, we extend TryOffDiff by adding class-specific embeddings to timestep embeddings [24].
-We demonstrate our results on the full DressCode dataset and show that the multi-garment TryOffDiff achieves comparable results to garment-specific variants.
-This makes TryOffDiff the first model to tackle multi-garment virtual try-off.
-Finally, we integrate TryOffDiff with state-of-the-art VTON models to enable person-to-person try-on.
-Our results show that this approach outperforms models capable of p2p-VTON by reducing unwanted attribute transfer, such as skin color, from the source person.
-To summarize, our primary contributions are:
-  • TryOffDiff: A diffusion model tuned for virtual try-off, excelling at garment reconstruction.
-  • Multi-Garment VTOFF: The first approach capable of handling multiple garment categories, and enabling simultaneous reconstruction of multiple garments, achieved through class-specific embeddings.
-  • Enhanced Person-to-Person VTON: Improved performance by feeding TryOffDiff output into VTON systems.
+In the rapidly evolving landscape of e-commerce, particularly within the fashion industry, the ability to provide customers with immersive and personalized shopping experiences is crucial for brands to differentiate themselves and drive engagement.
+Virtual Try-On (VTON) [19] technology has emerged as a powerful tool in this regard, allowing customers to visualize how garments would look on them without physically trying them on.
+Traditional VTON systems rely on standardized product images from e-commerce catalogs [8, 22], which may not fully capture the nuances of how a garment appears when worn by a real person, and are often unavailable for vintage, user‑generated, or unbranded items.
+These constraints can reduce the authenticity of the virtual try-on experience, potentially affecting customer satisfaction and conversion rates.
+Virtual Try‑Off (VTOFF) [32] overcomes these limitations by reconstructing standardized garment images directly from photos of people wearing them, addressing the scarcity of standardized product shots.
+Unlike VTON, which synthesizes a person wearing a specified garment, VTOFF extracts the garment itself in a clean, catalog-style format, free of stylistic variations.
+This capability is highly valuable in advertising and marketing, where visuals assets strongly influence purchasing behavior [36, 31].
+Traditional production of catalog imagery demands specialized equipment and extensive post‑processing, making it both time‑consuming and costly.
+By contrast, VTOFF enables rapid generation of consistent, high‑quality garment visuals without a dedicated studio setup, streamlining product promotion and boosting customer engagement.
+Furthermore, VTOFF supports visual retrieval from user-generated content, enabling automated trend analysis by isolating garments from influencer or street-style images, thus enriching brand intelligence and market research.
+Moreover, VTOFF opens new possibilities for person-to-person Virtual Try-On (p2p-VTON) [37], a more challenging variant of VTON that uses a photo of another person wearing the desired garment as the conditioning input.
+By integrating VTOFF with traditional VTON pipelines, p2p-VTON becomes feasible without requiring standardized catalog images.
+This functionality is especially appealing in social commerce, where peer influence drives purchasing decisions.
+By leveraging VTOFF, brands can create interactive, socially engaging shopping experiences, boosting conversion rates and fostering stronger connections with their audiences.
+In this paper, we present Multi-Garment TryOffDiff (MGT), a diffusion-based model designed to address the unique challenges of VTOFF task.
+Built upon Stable Diffusion architecture, MGT replaces text conditioning with image conditioning, utilizing SigLIP features [42] processed through an adapter module to capture garment-specific properties such as texture, shape and patterns.
+This approach yields competitive results on the VITON-HD dataset for upper-body garment reconstruction, despite not being trained on it, highlighting robust cross-domains generalization.
+To handle multi-garment scenarios, we extend TryOffDiff [32] by incorporating class-specific embeddings into the timestep embeddings [24].
+Evaluated on the full DressCode dataset, MGT performs comparably to garment-specific models, establishing it as the first model to support multi-garment VTOFF.
+Additionally, integrating MGT with state-of-the-art VTON models enhances p2p-VTON by minimizing unwanted attribute transfer, such as skin color, from the source person.
+Our key contributions are:
+  • Multi-Garment TryOffDiff: A diffusion-based virtual try-off model that introduces class-specific embeddings to enable simultaneous reconstruction of multiple garment categories, delivering robust cross-domain generalization and marking the first approach to support multi-garment try-off scenarios.
+  • Integration of VTOFF outputs with existing try-on pipelines, enabling p2p-VTON with improved preservation of person-specific attributes and reduced unwanted attribute transfer, resulting in higher-fidelity outputs.
 Figure 2: Overview of various fashion image generation pipelines.
@@
 2 Related Work
-This section reviews prior research relevant to our TryOffDiff model, covering Virtual Try-Off, Image-based Virtual Try-On, and Conditional Diffusion Models.
-Virtual Try-Off generates standardized garment images from clothed individuals, complementing virtual fitting tasks.
-Image-based Virtual Try-On synthesizes personalized garment images using catalog or person-worn inputs, supporting advanced variants.
-Conditional Diffusion Models offer the generative backbone for our approach, enabling controlled synthesis through tailored conditioning mechanisms.
-These areas highlight the challenges and strategies shaping our work on multi-garment reconstruction.
+This section reviews prior work on Virtual Try-Off, Image-based Virtual Try-On, and Conditional Diffusion Models.
+Virtual Try-Off generates standardized garment images from clothed individuals.
+Image-based Virtual Try-On synthesizes images of a person wearing a given garment.
+Conditional Diffusion Models serve as the generative backbone of recent approaches, supporting controlled synthesis through tailored conditioning mechanisms.
+These areas form the basis for our work on multi-garment reconstruction.
 Virtual Try-Off.
-Virtual Try-Off (VTOFF) generates standardized garment images from photos of clothed persons, a task with growing interest due to its applications in e-commerce.
-Early works, such as TileGAN [42], used a two-stage approach: a U-Net-inspired encoder-decoder to transfer features like color and shape for a coarse image, followed by a pix2pix-based refinement step.
-With the rise of general-purpose text-to-image diffusion models like Stable Diffusion [28], later works shifted toward text-driven garment generation.
-ARMANI [47], DiffCloth [48], GarmentAligner [46], and MGD [2] explored finetuning latent diffusion models on garment-caption datasets.
-However, text-based approaches require highly detailed captions, which are difficult to produce, and often struggle to follow instructions precisely, leading to suboptimal outputs.
-To address these limitations, recent approaches adopted image-based conditioning.
-IGR [31] finetuned Stable Diffusion v1.5 with a dual-tower architecture.
-Similarly, RAGDiffusion [32] employed SDXL [26] with a dual-tower design and a coarse mask from a complex retrieval-augmented generation.
-TryOffAnyone [37] finetuned Stable Diffusion v1.5 for VTOFF, training only self-attention layers (similar to CatVTON [10]), without additional image encoders.
-This resulted in a lightweight yet powerful model, making it a strong baseline.
-Despite these advances, VTOFF research faces challenges: many studies are not publicly available, limiting open exploration, and most focus solely on upper-body garments, leaving broader garment categories underexplored.
+Virtual Try-Off (VTOFF) [32] generates standardized garment images from photos of dressed people, a task with growing interest due to its applications in e-commerce.
+Early works, such as TileGAN [41], used a two-stage pipeline: a U-Net-like encoder-decoder for coarse synthesis followed by a pix2pix-based refinement.
+With the advent of text-to-image diffusion models such as Stable Diffusion [27], recent works explored text-guided garment generation.
+ARMANI [46], DiffCloth [47], GarmentAligner [45], and MGD [2] finetuned latent diffusion models using garment-caption datasets.
+However, text-based methods require detailed captions and often fail to follow them precisely, limiting quality.
+To address this, image-based conditioning was introduced.
+TryOffDiff [32] formalized the VTOFF task and replaced text embeddings in Stable Diffusion v1.4 with image embeddings.
+IGR [30] proposed a dual-tower variant of Stable Diffusion v1.5.
+TryOffAnyone [35] finetuned only self-attention layers, following CatVTON [10], enabling a lightweight model with competitive results.
+Despite progress, VTOFF remains underexplored.
+Many models are not publicly available, and most focus solely on upper-body garments, limiting generalizability.
 Image-based Virtual Try-On.
-The objective of image-based virtual try-on is to create composite images of a target person wearing a specific garment, preserving the person’s identity, pose, and body shape while accurately rendering garment details.
-CAGAN [19] introduced this task with a cycle-GAN approach, while VITON [15] formalized it as a two-step, supervised framework: warping the garment through non-parametric geometric transformations [4], followed by blending it onto the person.
-CP-VTON [35] improved warping with a learnable thin-plate spline (TPS) transformation via a geometric matcher, later enhanced by dense flow [16] and appearance flow [14] for better pixel-level alignment of textures and folds.
-Despite these advanves, warping-based methods struggled with complex garment features like logos and intricate patterns.
-Recent efforts shifted to GAN-based and diffusion-based models to overcome these limitations.
-FW-GAN [12] synthesized try-on videos, while PASTA-GAN [38] adapted StyleGAN2 for person-to-person try-on.
-However, GANs faced challenges with unstable training and mode collapse, prompting a move toward diffusion models, which offer greater stability.
-M&M-VTO [49] introduced a single-stage diffusion model for multi-garment try-on using one person image and multiple garment inputs.
-IDM-VTON [9] used two modules to encode garment semantics, extracting high- and low-level features with cross-attention and self-attention layers.
+Image-based virtual try-on (VTON) generates images of a person wearing a specific garment, preserving identity, pose, and body shape while accurately rendering garment details.
+CAGAN [19] introduced this with a cycle-GAN framework.
+VITON [15] formalized it as a two-step pipeline: warping the garment using non-parametric geometric transformations [4], then blending it with the person image.
+CP-VTON [33] introduced a learnable thin-plate spline (TPS) transformation via a geometric matcher, later enhanced by dense flow [16] and appearance flow [14] to better align textures and folds.
+However, warping-based methods struggle with occluded images as they lack generative capabilities.
+Recent efforts shifted to GANs and diffusion models.
+FW-GAN [12] targeted try-on videos.
+PASTA-GAN [37] applied StyleGAN2 to person-to-person try-on.
+GANs, however, suffer from instability and mode collapse, which diffusion models avoid.
+IDM-VTON [9] used two modules to encode garment semantics, extracting high- and low-level features with cross- and self-attention layers.
 OOTDiffusion [39] leveraged pretrained latent diffusion models, integrating garment features into a denoising UNet via outfitting fusion.
 In a lighter approach, CatVTON [10] avoided heavy feature extraction, proposing a compact model from a pretrained latent diffusion model with promising results using fewer parameters.
-Adapting VTON models for VTOFF is challenging, as VTON often relies on additional inputs like text prompts, keypoints, or segmentation masks, requiring careful adjustment.
-Importantly, while both VTON and VTOFF tasks involve garment manipulation, they differ fundamentally.
-VTON models have access to complete garment details, allowing them to primarily focus on warping the item to fit a target pose.
-In contrast, VTOFF models must work with only partial garment information from a reference image, where occlusions and deformations are common, requiring them to reconstruct missing details from limited visual cues.
+Adapting VTON models for VTOFF is ineffective, as VTON often relies on additional inputs such as text prompts, keypoints, or segmentation masks, requiring careful adjustment [32].
+More importantly, while both VTON and VTOFF tasks involve garment manipulation, they differ fundamentally.
+VTON has access to complete garment information and focuses on adapting it to a given pose.
+In contrast, VTOFF must recover standardized garments from partially visible, occluded, or deformed inputs, requiring the model to infer missing details.
 Conditional Diffusion Models.
-Latent Diffusion Models [28] (LDMs) achieved great success in recent years, providing control over generative process through cross-attention mechanisms [33].
-These models support conditioning on various input modalities, including text [5, 3, 13] and images [29, 25, 30].
-In text-guided synthesis, ControlNet [44] and T2I-Adapter [23] enhance pretrained models with additional layers for precise spatial control.
-IP-Adapter [41] further improves flexibility by decoupling cross-attention for text and image features, allowing image-guided generation with optional structural conditions.
-Prompt-Free Diffusion [39] eliminates text prompts entirely, relying on reference image and optional structural inputs to guide output.
-Despite progress, applying these models to garment reconstruction poses challenges.
-Text-guided methods require highly detailed prompts for each sample to define product attributes, which is impractical.
-Similarly, existing image-guided models lack the ability to enforce the strict standards of fashion photography [7], limiting their suitability.
+Latent Diffusion Models [27] (LDMs) generate high-quality images using cross-attention for conditioning on various modalities, including text [5, 3, 13] and images [28, 25, 29].
+Text-guided methods such as ControlNet [43] and T2I-Adapter [23] improve spatial control with auxiliary networks.
+IP-Adapter [40] further improves flexibility by decoupling cross-attention for text and image features, allowing image-guided generation.
+Prompt-Free Diffusion [38] eliminates text prompts entirely, relying on reference image and optional structural inputs for guidance.
+While effective for general image synthesis, existing conditional diffusion models are not well suited for garment reconstruction.
+Text-guided methods require highly detailed and consistent prompts to specify product attributes, which is labor-intensive and difficult to scale.
+Image-guided approaches often lack the precision required for fashion photography [7].
+As shown in TryOffDiff [32], applying these models directly to VTOFF results in low-fidelity outputs.
 3 Methodology
-Figure 3: Overview of TryOffDiff.
+Figure 3: Overview of MGT.
 Given a reference image and a class label (e.g.
@@
 The model, with trainable Adapter, Embedding, and U-Net components, produces a standardized garment image, which is decoded by a frozen VAE.
-This section formally defines the virtual try-off task and outlines TryOffDiff model, including SigLIP-based image conditioning and class-conditioning for multi-garment support.
+This section formally defines the virtual try-off task and outlines Multi-Garment TryOffDiff model.
 3.1 Virtual Try-Off
-Problem Formulation.
-Consider an RGB image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ with height $H\in\mathbb{N}$ and width $W\in\mathbb{N}$, representing a person wearing garments.
-The VTOFF task is to produce a standardized product image $\mathbf{G}\in\{0,\ldots,255\}^{H\times W\times 3}$, that meets commercial catalog standards.
+Consider an RGB image $\mathbf{I}\in\{0,\ldots,255\}^{H\times W\times 3}$ with height and width $H,W\in\mathbb{N}$, representing a person wearing garments.
+The VTOFF task aims to produce a standardized product image $\mathbf{G}\in\{0,\ldots,255\}^{H\times W\times 3}$, that meets commercial catalog specifications.
 Formally, the goal is to train a generative model that learns the conditional distribution $P(G|C)$, where $G$ and $C$ represent the variables corresponding to garment images and reference images (serving as condition), respectively.
@@
 Then, given a specific reference image $\mathbf{I}$ as conditioning input, the objective is for a sample $\hat{\mathbf{G}}\sim Q(G|C=\mathbf{I})$ to resemble a true sample of a garment image $\mathbf{G}\sim P(G|C=\mathbf{I})$ as closely as possible.
-3.2 TryOffDiff
-The architecture of TryOffDiff mirrors Stable Diffusion [28] (v1.4), a latent diffusion model originally designed for text-to-image generation using CLIP’s [27] text encoder.
-We replace text prompts for direct image-guided generation.
+3.2 Multi-Garment TryOffDiff (MGT)
+The product image $\mathbf{G}$ may contain various types of garments, such as upper-body wear, lower-body wear, or full-body dresses.
+Reconstructing a specific garment type from full-body photos makes the virtual try-off task more challenging and necessitates additional guidance.
+To address this, we extend the existing VTOFF model TryOffDiff [32] with class-specific embeddings that help disentangle and reconstruct individual garment categories.
 Image Conditioning.
-A core challenge in image-guided generation is effectively incorporating visual features into the conditioning mechanism of the generative model.
-CLIP’s ViT [27] has become a popular choice for image feature extraction due to its general-purpose capabilities.
-Recently, SigLIP [43] introduced modifications that improve performance, particularly for tasks requiring more detailed and domain-specific visual representations.
-Therefore, we use the SigLIP model as image feature extractor and retain the full sequence of token representations in its final layer to preserve spatial information essential for accurate garment reconstruction.
-These features are processed by a trainable adapter module:
+A central challenge in image-guided generation is effectively incorporating visual features into the generative model’s conditioning mechanism.
+CLIP’s ViT [26] is a common choice for image encoding due to its general-purpose representations, but recent improvements by SigLIP [42] make it better suited for detailed visual tasks.
+We adopt SigLIP as the image encoder and extract the full sequence of tokens from its final layer to retain spatial information.
+These token embeddings are projected and normalized (LN) [1] via a lightweight adapter module:
 $\mathbf{C}(\mathbf{I})=(\mathrm{LN}\circ\mathrm{Linear}\circ\mathrm{SigLIP})(\mathbf{I})\in\mathbb{R}^{n\times m}$ (1)
-where SigLIP embeddings are projected by a linear layer and normalized (LN) [1], following a structure similar to IP-Adapter [41].
-The adapted features $\mathbf{C}(\mathbf{I})$ are then incorporated into the denoising U-Net’s cross-attention layers.
+where $\mathbf{C}(\mathbf{I})$ denotes the adapted feature sequence, and $n,m\in\mathbb{N}$ denote the token and feature dimensions, respectively.
+This design follows the IP-Adapter [40].
+The resulting features are injected into the cross-attention layers of the denoising U-Net.
 Specifically, the key $\mathbf{K}$ and value $\mathbf{V}$ of the attention mechanism at each layer are derived from the image features through linear transformations:
@@
 where $\mathbf{W}_{k}\in\mathbb{R}^{m\times d_{k}}$ and $\mathbf{W}_{v}\in\mathbb{R}^{m\times d_{v}}$.
-This mechanism allows the U-Net to use the reference image’s visual details, improving the precision of the generated garment.
-[…]
-[…]
-TryOffAnyone* [37] - - - 72.0 - - 17.2 25.3 - 8.9 21.0
-TryOffDiff-single CLIP ViT-B/32 - (257,1024) - 77.9 69.2 46.4 35.3 16.8 7.7 5.2 23.7
-CLIP ViT-B/32 Linear+LN (77,768) - 78.8 70.5 46.5 34.3 16.3 7.4 4.7 23.5
-SigLIP-B/16 Linear+LN (77,768) - 78.4 69.8 47.3 33.1 16.4 6.6 5.1 22.1
-CLIP ViT-B/32 Linear+LN (77,768) 79.8 71.5 47.8 32.0 15.2 5.8 4.1 21.5
-SigLIP-B/16 Linear+LN (77,768) 80.3 72.5 49.2 30.1 14.5 5.7 3.9 20.3
-Table 1: Training configurations and quantitative results of baseline and ablation studies on VITON-HD-test.
-The “Pretrain” check mark indicates whether the U-Net is initialized from Stable Diffusion v1.4 (✓) or trained from scratch (-).
-All ablations mirror the model architecture of Stable Diffusion (v1.4).
-*Values for TryOffAnyone are taken from the original paper.
+This conditioning has already proven effective for VTOFF results, as demonstrated by TryOffDiff [32].
 Garment Type Conditioning.
-Reference images often contain multiple garments, such as a shirt and pants, requiring the model to distinguish which garment to reconstruct.
-To resolve this ambiguity and enable multi-garment support, we condition the model on text embeddings corresponding to garment categories—“lower body”, “upper body”, or “dress”—aligned with the class labels of the Dress Code dataset.
-These text embeddings are passed through a linear layer and then combined with timestep embeddings [24] via element-wise addition before being fed into the residual blocks of the diffusion model.
-This conditioning mechanism guides the denoising process to generate the garment corresponding to specified category.
+Multi-Garment TryOffDiff (MGT) extends TryOffDiff [32] by introducing a simple class conditioning mechanism to support multiple garment types within a single model.
+Reference images often contain several garments (e.g., shirts and pants), requiring the model to distinguish which garment to reconstruct.
+To resolve this ambiguity and enable multi-garment support, we condition the model on garment category labels—“lower body”, “upper body”, or “dress”—consistent with the class annotations in the DressCode dataset.
+Each category is associated with a learnable embedding vector.
+Let $\mathcal{C}=\{1,2,3\}$ denote the index set of garment classes and $\mathbf{E_{c}}\in\mathbb{R}^{|\mathcal{C}|\times d}$ be the learnable embedding matrix, where $d=1280$ matches the diffusion model’s timestep embedding dimension.
+Given a class index $c\in\mathcal{C}$, the corresponding garment type embedding is retrieved as:
+$\mathbf{e}_{c}=\mathbf{E_{c}}[c]\in\mathbb{R}^{d}$ (3)
+Let $\mathcal{T}=\{1,\cdots,T\}$, where $T=1000$, denote the index set of timesteps and $\mathbf{E_{t}}\in\mathbb{R}^{T\times d}$ be the learnable timestep embedding matrix.
+Given a timestep index $t\in\mathcal{T}$, the corresponding timestep embedding is retrieved as:
+$\mathbf{e}_{t}=\mathbf{E_{t}}[t]\in\mathbb{R}^{d}$ (4)
+Then, the conditioned timestep embedding is defined as the element-wise addition of the timestep embedding and the garment type embedding:
+$\mathbf{e_{t}}_{\text{cond}}=\mathbf{e}_{t}+\mathbf{e}_{c}$ (5)
+This conditioned timestep embedding is injected into each residual block of the denoising U-Net (following [24]), where it modulates feature activations to guide generation towards the target garment class.
+This mechanism allows the model to disambiguate between garments types and produce semantically accurate outputs.
 4 Experiments
-Since VTOFF task is newly introduced, few open-source methods exist.
-We evaluate our proposed TryOffDiff model against the publicly available TryOffAnyone [37] to establish a baseline.
-To ensure reproducibility, we provide a detailed description of our experimental setup.
-We adopt DISTS [11] as the primary evaluation metric, supplemented by standard metrics for a comprehensive comparison.
-Qualitative results are also presented to showcase TryOffDiff’s ability to handle diverse and challenging inputs.
+Since Virtual Try-Off is a newly introduced task, few open‐source methods exist.
+We evaluate our Multi-Garment TryOffDiff (MGT) against two publicly-available baselines: TryOffDiff [32] and TryOffAnyone [35], to establish benchmarks.
+We adopt DISTS [11] as our primary metric and report additional standard measures for a comprehensive evaluation.
+We also present qualitative examples to illustrate MGT’s performance on diverse inputs.
 4.1 Experimental Setup
@@
 DressCode contains $53,792$ high-resolution ($1024\times 768$) person-garment pairs, with $48,392$ pairs for training–comprising dresses ($27,678$), upper-body ($13,563$), and lower-body ($7,151$)–and $5,400$ pairs for testing, evenly split with $1,800$ pairs per category.
-For person-to-person try-on, where ground-truth images are unavailable, we randomly pair garments across individuals and provide a text file of input image filenames for reproducibility.
+For person-to-person (p2p) try-on, where ground-truth are unavailable, we randomly pair garments across individuals and provide a text file of input image filenames for reproducibility.
 For DressCode, a label (upper-body, lower-body, or dresses) indicates the garment to transfer; this is not needed for VITON-HD, which includes only upper-body garments.
@@
 Implementation Details.
-Denoising U-Net weights are initialized from Stable Diffusion v1.4 [28] and subsequently fine-tuned.
+Denoising U-Net weights are initialized from Stable Diffusion v1.4 [27] and subsequently finetuned.
 The Adapter module and class Embedding layer are trained from scratch.
@@
 On a single NVIDIA A40 GPU, inference on DressCode test set takes 2.8 seconds per image and requires 10.1GB of memory.
-4.2 Ablation Studies
-We perform ablation studies to assess key components of TryOffDiff, including the image encoder, adapter module, conditioning feature dimensions, and the impact of finetuning a pretrained U-Net versus training from scratch.
-The configurations for these experiments are summarized in Tab.
-1.
-In addition, Fig.
-4 illustrates the impact of guidance scale and inference steps on FID and DISTS scores, evaluated on the Dress Code-test set using the Euler noise scheduler.
-This experiment reveals how these hyperparameters influence reconstruction quality, providing insights into optimal inference setting for TryOffDiff.
+4.2 Quantitative Results
+We compare MGT to TryOffDiff [32] and TryOffAnyone [35] on the VITON-HD and DressCode datasets, with results reported in Table 1 (DressCode), Table 2 (VITON-HD), and Table 3 (p2p-VTON).
+TryOffDiff and TryOffAnyone are trained on VITON-HD (upper-body only), while MGT is trained on DressCode, supporting multiple garment categories (upper-body, lower-body, dresses).
+[…]
+Upper-body
+TryOffDiff-single 80.8 73.8 47.8 31.6 17.1 5.2 4.7 21.6
+MGT (ours) 80.2 73.2 48.1 32.3 17.8 5.3 5.5 22.2
+Lower-body
+TryOffDiff-single 81.1 76.8 55.5 30.0 22.6 9.4 5.9 21.1
+MGT (ours) 80.4 75.6 54.1 30.8 22.4 9.0 6.8 21.7
+Dresses
+TryOffDiff-single 81.6 76.1 55.6 26.1 18.4 8.2 4.4 20.8
+MGT (ours) 81.6 76.1 55.5 26.4 18.9 8.0 4.6 21.2
+Table 1: Quantitative results on DressCode-test. Comparing TryOffDiff(single-category) to MGT(unified model).
+No prior VTOFF model handles multiple garment categories.
+Therefore, to establish a baseline and to measure the cost of unified modeling, we train three category-specific TryOffDiff variants on DressCode (upper-body, lower-body, dresses) and compare them against MGT, a single model conditioned on garment type.
+This setup tests whether supporting multiple categories in a unified architecture results in performance degradation.
+Table 1 shows MGT achieves comparable performance across all categories, with minor trade-offs (e.g., DISTS: 22.2 vs.
+21.6 for upper-body, LPIPS: 30.8 vs.
+30.0 for lower-body, and FID: 18.9 vs.
+18.4 for dresses).
+This shows that a single, unified approach can replace multiple specialized models without significant performance loss.
+[…]
+TryOffDiff [32] 79.5 70.4 46.2 32.4 25.1 9.4 8.9 23.0
+TryOffAnyone [35] 71.9 - - 17.2 25.3 - 2.0 21.0
+[…]
+[…]
+We further assess MGT’s generalization in a cross-dataset setting by evaluating it on VITON-HD, despite being trained exclusively on DressCode.
+As shown in Table 2, MGT outperforms TryOffDiff in FID (21.9 vs.
+25.1) and remains competitive across other metrics.
+This indicates strong generalization despite the domain shift.
+VITON-HD DressCode
+[…]
+CatVTON [10] 12.0 3.5 3.9 8.4 1.9 3.1
+OOTD [39] + Ground truth 10.8 2.8 2.0 7.5 3.4 2.5
+OOTD [39] + MGT (ours) 11.9 3.3 2.6 7.9 3.4 2.7
+Table 3: Quantitative results for p2p-VTON.
+We evaluate OOTDiffusion (OOTD) with ground truth (GT) garments and MGT-predicted garments, alongside CatVTON, a specialized p2p-VTON model.
+Our model’s output, when integrated with a VTON model, achieves competitive performance, even though it is not explicitly trained for the p2p task.
+In p2p-VTON setup (Tab.
+3), we pair MGT with OOTDiffusion and compare the results to CatVTON, a dedicated p2p-VTON model.
+On VITON-HD, OOTD+MGT slightly outperforms CatVTON.
+On DressCode, MGT achieves competitive results, especially in perceptual metrics, showing its effectiveness in tasks beyond its training objective.
+These findings highlight MGT’s versatility across both VTOFF and p2p-VTON tasks.
+Its improvements in perceptual quality and structural fidelity translate to practical benefits in realistic garment rendering.
+Lastly, Figure 4 explores the influence of inference hyperparameters (guidance scale and number of steps) on FID and DISTS scores.
+This analysis provides insights into optimal inference configurations for MGT.
 (a) Guidance Scale
 (b) Inference steps
-Figure 4: Impact of guidance scale ($s$) and inference steps ($n$) on DISTS and FID scores. Experiments are conducted on DressCode-test with TryOffDiff using the Euler scheduler [20].
-4.3 Quantitative Results
-We evaluate TryOffDiff’s performance on the VITON-HD and Dress Code datasets, comparing it to baseline TryOffAnyone [37] and category-specific ablations.
-Results are presented in Tab.
-1 (VITON-HD), Tab.
-2 (Dress Code), and Tab.
-3 (p2p-VTON).
-[…]
-Dresses
-TryOffDiff-single 81.6 76.1 55.6 26.1 18.4 8.2 4.4 20.8
-TryOffDiff-multi 81.6 76.1 55.5 26.4 18.9 8.0 4.6 21.2
-Upper-body
-TryOffDiff-single 80.8 73.8 47.8 31.6 17.1 5.2 4.7 21.6
-TryOffDiff-multi 80.2 73.2 48.1 32.3 17.8 5.3 5.5 22.2
-Lower-body
-TryOffDiff-single 81.1 76.8 55.5 30.0 22.6 9.4 5.9 21.1
-TryOffDiff-multi 80.4 75.6 54.1 30.8 22.4 9.0 6.8 21.7
-All
-TryOffDiff-merged 81.2 75.6 52.9 29.2 11.4 6.7 3.1 21.2
-TryOffDiff-multi 78.6 75.4 57.4 25.0 12.8 6.8 4.1 21.3
-Table 2: Quantitative comparison of multi-garment VTOFF on DressCode-test.
-We compare category-specific models (TryOffDiff-single) trained independently on dresses, upper-body, and lower-body garments against a class-conditioned model (TryOffDiff-multi) trained jointly across all garment types.
-Additionally, we report results for TryOffDiff-merged, which aggregates the outputs of the three category-specific models into a single set.
-While non-distributional metrics (e.g., SSIM, LPIPS) for merged results reflect simple averages across categories, distributional metrics (e.g., FID, KID) are computed over the merged output set, capturing holistic distributional properties.
-VITON-HD Dress Code
-[…]
-CatVTON 12.0 3.5 3.9 8.4 1.9 3.1
-OOTD + GT 10.8 2.8 2.0 7.5 3.4 2.5
-OOTD + TryOffDiff 11.9 3.3 2.6 7.9 3.4 2.7
-Table 3: Quantitative results for p2p-VTON.
-We evaluate OOTDiffusion (OOTD) with ground truth (GT) garments and TryOffDiff-predicted garments, alongside CatVTON, a specialized p2p-VTON model.
-Our TryOffDiff model, when integrated with a VTON model, achieves competitive performance, even though it is not explicitly trained for the p2p task.
-The results in Tab.
-1 reveal key insights into our design choices.
-First, finetuning a pretrained U-Net consistently outperforms training from scratch across all metrics, underscoring the advantage of leveraging pretrained weights for initializing the diffusion model.
-Second, the SigLIP encoder outperforms CLIP, especially in perceptual quality metrics like FID and DISTS, suggesting that SigLIP’s visual representations are more effective for detailed garment reconstruction.
-Finally, incorporating the adapter module further improves performance by more effectively conditioning the diffusion process with image features.
-On VITON-HD dataset (Tab.
-1), our method outperforms the baseline across all metrics except LPIPS, where TryOffAnyone achieves a better score, though TryOffDiff excels in global metrics like FID and DISTS which is our primary metric.
-In our evaluation on the Dress Code dataset (Tab.
-2), we consider: (i) category-specific TryOffDiff models trained individually on dresses, upper-body, and lower-body garments; (ii) a class-conditioned TryOffDiff model trained jointly on all categories using garment-type embeddings; and (iii) an aggregated result, denoted TryOffDiff-merged, where outputs from the three category-specific models are combined post hoc.
-The class-conditioned model remains competitive across all categories, demonstrating its robustness as a unified solution.
+Figure 4: Impact of guidance scale ($s$) and inference steps ($n$) on DISTS and FID scores. Evaluated on DressCode-test with MGT using the Euler scheduler [20].
 Figure 5: Qualitative comparison of garment-specific and Multi-Garment TryOffDiff.
-First row displays the conditioning image.
-Second row shows dataset-specific reconstructions produced by TryOffDiff model trained on individual datasets: VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12).
+First row displays the reference image.
+Second row shows dataset-specific reconstructions produced by TryOffDiff model trained on single category: VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12).
 Last row shows predictions of Multi-Garment TryOffDiff trained on full DressCode with garment type conditioning.
-For person-to-person virtual try-on (Tab.
-3), we integrate TryOffDiff with OOTDiffusion and compare it to CatVTON, a specialized p2p-VTON model.
-On VITON-HD, OOTD+TryOffDiff performs slightly better than CatVTON.
-On Dress Code, TryOffDiff, when paired with OOTD, delivers competitive performance, approaching specialized p2p-VTON models like CatVTON, especially in perceptual quality metrics.
-These findings highlight TryOffDiff’s versatility and effectiveness, enabling high-quality garment reconstruction for both VTOFF and p2p-VTON tasks.
-Its improvements in structural similarity and perceptual quality translate into practical benefits, such as more realistic garment rendering in both VTOFF and p2p-VTON applications.
+(a) Input
+(b) ToD [32]
+(c) ToA [35]
+(d) MGT
+(e) GT
+Figure 6: Qualitative comparison on VITON-HD.
+Despite not being trained on VITON-HD, MGT achieves competitive results compared to baselines specifically trained on it: TryOffDiff (ToD) [32] and TryOffAnyone (ToA) [35].
 4.3 Qualitative Analysis
-Figure 5 illustrates the performance of TryOffDiff in garment reconstruction, comparing two variants: garment-specific models (row 2) and multi-garment model (row 3).
-The conditioning images, displayed in first row, are sourced from VITON-HD (cols 1-3), DressCode upper-body (cols 4-6), lower-body (cols 7-9), and dresses (cols 10-12).
-Garment-specific models excel in precision, particularly for text and fine details (cols 2 and 5), likely due to the pretrained model initialization.
-In contrast, the multi-garment model offers versatility, delivering strong performance across diverse garment types and datasets, albeit with slight trade-offs in detail compared to specialized models.
-Nevertheless, it reconstructs a high level of detail (column 6) and shows robust generalization to VITON-HD, which was unseen during training.
-In some cases (cols 2, 7 and 9), it even produces better fidelity of color.
-Figure 6 illustrates the results of p2p-VTON.
-Overall, neither method stands out as a definitive winner.
-CatVTON excels in preserving texture and pattern details but occasionally introduces diffusion artifacts (row 2).
-It can also transfer unintended attributes, such as skin color, from target model to the source model (row 1)–a limitation not observed in classical try-on models.
-OOTDiffusion combined with TryOffDiff avoids such attribute transfer issues but may struggle due to masking, resulting in residual source garment attributes in the output (rows 3 and 4).
-These results highlight the potential of Virtual Try-Off task and the TryOffDiff model.
-Although TryOffDiff was not specifically trained for p2p-VTON, its integration with VTON models presents a promising approach, already demonstrating competitive performance compared to state-of-the-art person-to-person virtual try-on methods.
-Figure 6: Qualitative comparison for Person-to-Person Virtual Try-On.
-Columns show: (a) person to be dressed, (b) CatVTON output using a person image with target garment as condition for direct p2p-VTON, (c) OOTDiffusion output with ground truth target garment, and (d) OOTDiffusion output with multi-garment TryOffDiff prediction.
+Figure 5 shows outputs from two sets of models: category-specific TryOffDiff variants (row 2) and our unified MGT model (row 3).
+Top row shows input images drawn from; VITON-HD (cols 1–3), DressCode upper-body (4–6), lower-body (7–9), and dresses (10–12).
+While all models produce visually plausible outputs, MGT matches or surpasses single-category models in several cases.
+It maintains clear text (cols 1, 6), emblems (col 4), and accurate garment shapes (col 5), and recovers detailed textures (cols 2, 10).
+MGT also performs strongly on the unseen VITON-HD dataset (cols 1–3), demonstrating robust generalization across garment types and domains.
+To complement the quantitative results on the VITON-HD (Tab.
+2), we provide a qualitative comparison in Figure 6.
+Despite not being trained on VITON-HD, MGT produces outputs that are visually competitive with baselines that were trained specifically on this dataset, including TryOffDiff [32] and TryOffAnyone [35].
+It maintains structural coherence and texture detail, even under challenging conditions involving complex patterns or non-frontal poses.
+These findings affirm MGT’s ability to generalize without the need for per-category or per-dataset tuning.
+Figure 7 shows results for p2p-VTON.
+CatVTON transfers textures and patterns accurately but occasionally introduces artifacts (rows 1–2) or unintended changes such as skin tone shifts (row 1).
+In contrast, combining OOTDiffusion with MGT yields consistent, artifact-free outputs.
+Although no method dominates in all scenarios, the decoupling of garment reconstruction (via MGT) from rendering (via VTON) results in more interpretable and stable outputs.
+(a) Single Garment
+(b) Multi-Garment
+Figure 7: Qualitative comparison for Person-to-Person Virtual Try-On in single and multi-garment settings.
+Columns: (i) source model image (person to be dressed), (ii) output from CatVTON, using a person image with target garments as condition for direct p2p-VTON, (iii) output from OOTDiffusion (OOTD), conditioned on the ground-truth (GT) target garments, and (iv) output from OOTDiffusion with our Multi-Garment TryOffDiff (MGT) pipeline.
+Each result shows the generated outfit on the source model, with the corresponding target garments shown in the top-left and top-right corners.
+For single-garment settings (a), only the top-left garment is applied.
+For multi-garment settings (b), garments are applied sequentially: the top-left garment is applied first, and its output serves as input for applying the top-right garment.
 5 Conclusion
-In this paper we propose TryOffDiff, a latent diffusion model tailored to Virtual Try-Off.
-We substitute text conditioning mechanism with adapted SigLIP features to steer the generative process, enabling effective garment reconstructions.
-TryOffDiff achieves state-of-the art results on VITON-HD.
-Further experiments were conducted on DressCode, which contains three types of garments: upper-body, lower-body and dresses.
-While the baseline architecture can be trained separately for each category, we enhance TryOffDiff by introducing class conditioning.
-This advancement allows a single model to perform multi-garment reconstruction, successfully capturing all garment types with minimal loss in detail quality.
-To our knowledge, this is the first method to address multi-garment VTOFF in a unified framework.
-Moreover, we demonstrate that integrating VTOFF with VTON models enables Person-to-Person Virtual Try-On, achieving performance comparable to specialized models.
-This pipeline offers practical advantages: it leverages existing VTON models for a wider applications and decouples garment reconstruction from other generative artifacts, simplifying the assessment of p2p-VTON outcomes.
-Despite these promising results, challenges persist in preserving complex structures, such as logos and printed designs.
-Future work could explore advanced generative models, alternative visual conditioning methods and additional losses to enhance detail preservation.
+We introduced Multi-Garment TryOffDiff (MGT), a unified latent diffusion model for the Virtual Try-Off (VTOFF) task.
+By incorporating a lightweight class-conditioning mechanism, MGT supports the reconstruction of diverse garment types—upper-body, lower-body, and dresses—from a single reference image.
+Unlike prior work, which relies on separate models for each garment category, MGT achieves near-specialized performance within a single framework.
+On the DressCode dataset, it matches or surpasses category-specific TryOffDiff variants while supporting multiple categories simultaneously.
+To our knowledge, MGT is the first unified model to address multi-garment VTOFF.
+Moreover, MGT demonstrates strong cross-dataset generalization.
+Despite being trained exclusively on DressCode, it achieves competitive results on VITON-HD, highlighting its robustness to domain shifts and unseen garments.
+Beyond VTOFF, we show that MGT can be seamlessly integrated into person-to-person Virtual Try-On (p2p-VTON) pipelines.
+When paired with OOTDiffusion, it achieves performance on par with specialized p2p models like CatVTON, while offering a modular and interpretable pipeline, separating garment reconstruction from person rendering.
+Limitations and Future Work.
+MGT currently supports only three garment categories and does not handle layered clothing, constrained by the available dataset annotations.
+Fine-grained texture recovery and logo preservation also remain challenging, particularly for complex garments.
+Future research could focus on expanding garment coverage, refining conditioning mechanisms, and exploring higher-capacity architectures or perceptual training objectives to improve detail fidelity.
+In summary, MGT marks a step toward general-purpose virtual try-off by unifying garment reconstruction across categories in a single, adaptable model.
+Its strong performance, generalization capabilities, and modularity make it a valuable component for future virtual fashion systems.
 Acknowledgment